determine the Current in a long straight wire that would produce a magnetic field that produces a Magnetic Field four times that of Earth's Magnetic Field located at a distance 9.0 cm from the wire. (assuming that the Earth's Magnetic Field is equal to 5x10^(-5)T and \mu _(0)=4\pi x10^(-7)T. m/A).


1
Expert's answer
2022-05-22T14:38:59-0400
"B=\frac{\mu_0 I}{2\pi r}"

"I=\frac{2\pi r B}{\mu_0}=\frac{2\pi * 0.09*4*5*10^{-5}}{4\pi*10^{-7}}=90\:\rm A"